OpenStudy (anonymous):

find the inverse of the function and determine whether the inverse is a function: f(x)=7x^2

hero (hero):

We know \(y = f(x)\) so: \(y = 7x^2\) Now swap \(x\) and \(y\): \(x = 7y^2\) Next isolate y: \(y = \pm\sqrt{\frac{x}{7}}\) Replace \(y\) with \(f^{-1}(x)\) \(f^{-1}(x) = \pm\sqrt{\frac{x}{7}}\) The inverse doesn't appear to be a function since for every x value you get two y values. For a function, every x must produce exactly one y.

OpenStudy (anonymous):

thankyou!

hero (hero):

Hang on...I was posting a solution and then it cut out.